OpenStudy (anonymous):

sorry i wrote it wrong (6.2*10^7)/4.15*10^3)

OpenStudy (lalaly):

14939.8

OpenStudy (anonymous):

Divide the parts that aren't 10 to a power first... 6.2/4.15 = 1.49 (approx.) Then subtract the exponents on the powers of 10. 10^7/10^3 = 10^(7-3) = 10^4 So you get 1.49 x 10^4

OpenStudy (anonymous):

14939.8

OpenStudy (anonymous):

ty all:)

OpenStudy (anonymous):

I assumed you needed it still in scientific notation :)